OpenStudy (anonymous):

WILL GIVE MEDAL!!!!Expand (a – 1)(a – 2) by using the FOIL method.

OpenStudy (texaschic101):

(a - 1)(a - 2) F - multiply first terms in each set ....a * a = a^2 O - multiply outside terms ....a * 2 = 2a I - multiply inside terms .....-1 * a = -1a L - multiply last terms .....-1 * -2 = 2 Now put it all together a^2 + 2a - 1a + 2 now combine like terms a^2 + a + 2

OpenStudy (texaschic101):

any questions ?